# When Respondents Don’t Care Anymore: Identifying the Onset of Careless Responding

### v3

@@ Introduction
We propose to measure response invariability through a novel algorithm inspired by the longstring index of Johnson, (2005) that we call LongStringPattern (LSP). This algorithm exploits that invariable careless responses are characterized by content-independent response patterns. A pattern sequence has the defining property that it consists of repeated occurrences of the same response pattern. Consider the sequence of responses $-2-1-2-\ldots-1-$ 2 - … - 1 - 2, which consists of repeated occurrences of the pattern $1-2$. Denoting the length of a pattern by $l$, the above example exhibits a pattern of length $l=2$. Analogously, the sequence $-3-1-2-3-\ldots-1-2-$ 3 - … - 1 - 2 - 3 consists of repeated occurrences of the pattern $1-2-3$ of length $l=3$, whereas the straightlining sequence $1-1-\ldots-$ 1 - … - 1 consists of repeated occurrences of 1, which is a pattern of length $l=1$.
Define $l\textsf{-pattern}_{ij}$ as the number of consecutive items in the longest sequence of a recurring pattern of length $l$ in which the response of participant $i$ to item $j$ takes part, or assign the value 1 if the response is not part of a recurring pattern of length $l$. Consider the following two illustrative examples. First, for $l=2$, the response sequence $1-2-1-2-1-4-3-5-4-5-4$ will be assigned the 2-pattern sequence $4-4-4-4-1-1-1-4-4-4-4$ because the first four responses and last four responses comprise two occurrences of the distinct patterns $1-2$ and $5-4$, respectively, which are both of length $l=2$. The central three responses ($1-4-3$) are not part of any recurring pattern of length $l=2$, hence they are each assigned the value 1. Second, for a recurring pattern of length $l=1$ (i.e., consecutive identical responses), the response sequence $3-2-3-3-1-4-1-1-1$ is assigned the 1-pattern sequence $1-1-2-2-1-1-3-3-3$, since the subsequences of consecutive identical responses $3-3$ and $1-1-$ 1 - 1 comprise two and three responses, respectively.33endnote: 3The 1-pattern sequence can be seen as a generalization of the the longstring index of Johnson, (2005), which we recover by calculating $\max_{1\leq j\leq p}1\textsf{-pattern}_{ij}$ for each respondent $i=1,\dots,n$. In short, higher values in the $l$-pattern sequence capture a tendency to respond in fixed patterns of length $l$. If a participant starts engaging in invariable careless responding, we expect a changepoint in the $l$-pattern sequence (for some $l$) from relatively low to relatively high values, assuming that the survey design precludes such patterns in attentive responses.
@@
for participant $i=1,\dots,n$ and item $j=1,\dots,p$. We call the series of within-participant values $\textsf{LSP}_{ij}$ across all items a LongStringPattern (LSP) sequence. In this paper, we vary $l$ from 1 to the number of answer categories in the survey. The rationale behind this choice is that a careless respondent may repeat, e.g., the pattern $1-2-3-4-5$ in case of five answer categories. At the same time, we consider it unlikely that careless respondents choose complicated individual patterns whose length exceeds the number of answer categories.
We provide a detailed description of the LSP algorithm in Appendix B.3. An example of an LSP sequence is shown in Figure 3, which shows simulated responses of a hypothetical participant in a 300-item questionnaire (5 answer categories per item) who starts responding carelessly using the recurring pattern $3-4$ from the 260th item, resulting in systematically larger LSP values from that item onward.
@@
Note. The simulated participant started responding carelessly with the recurring pattern $3-4$ from the 260th item onward.
@@
As motivated in the previous two sections, we expect a changepoint in any of the measurements of carelessness from the onset of careless responding (cf. Figures 2 and 3), and no changepoint in the absence of carelessness. We therefore propose to apply the nonparametric cumulative sum self-normalization test of Shao & Zhang, (2010), which was developed in the context of time series analysis. It searches for the location of a possible changepoint in a multivariate series and has attractive theoretical guarantees (see Shao & Zhang,, 2010; Zhao et al.,, 2022, and our detailed description in Appendix B.4).
For each point in a multivariate series (we use measurements of carelessness associated with each questionnaire item), this changepoint detection test calculates a statistic that compares a quantity of interest (we take the mean) in the periods before and after the current point (item). If the maximum of this statistic exceeds a critical value, the point (item) where this maximum occurs is flagged as a changepoint. Otherwise no changepoint is flagged. The critical value thereby depends only on a pre-specified significance level $\alpha$ and the dimension of the multivariate series. Importantly, it controls the size of the test regardless of the distribution of the individual series. This gives CODERS great flexibility, as it allows us to use any combination of measurements of carelessness without making possibly unrealistic assumptions on their distributions or obtaining cutoff values in an ad-hoc manner. In line with literature that suggests that there should be overwhelming evidence when labeling respondents as careless (e.g., Huang et al.,, 2012), we recommend using a conservative significance level no greater than 1% for identifying changepoints.
As an illustrative example, consider Figure 4(a). It illustrates two dimensions computed from a given participant’s responses, namely autoencoder reconstruction errors and the LSP sequence. The former series has been shown before in Figure 2. The bottom panel of Figure 4(a) shows the test statistics computed from this two-dimensional series, where the maximum value at the 172nd item clearly exceeds the critical value. That is, CODERS identifies inconsistent carelessness from the 172nd item onward, as the reconstruction errors are in general much larger than before. Indeed, this simulated participant started choosing responses at random at the identified item. Figure 4(b) is similar, but computed from a different respondent, whose LSP sequence has been shown before in Figure 3. CODERS flags a changepoint at the 261st item. From this item onward, the LSP values are clearly elevated, indicating invariable carelessness. In fact, this simulated participant started repeating the pattern $3-4$ just one item before the estimated onset.
@@
Note. The left subfigure shows the output from CODERS for a simulated partially inconsistent respondent who starts responding randomly from the 172nd item onward, while the right subfigure does so for a simulated partially invariable respondent who starts repeating the pattern $3-4$ from the 260th item onward. Each subfigure visualizes the autoencoder reconstruction errors (RE), LongStringPattern (LSP) sequence, and test statistics associated with the corresponding participant’s responses. The blue solid vertical lines indicate the carelessness onset items identified by CODERS at significance level 0.1%, which correspond to the true onset item 172 for the partially inconsistent respondent, and to item 261 (one item after true onset) for the partially invariable respondent. The dashed horizontal line is the critical value of the test statistic at the 0.1% significance level (192.5; see Table B.2 in Appendix B.4). The reconstruction errors in the left subfigure are those shown in Figure 2, and the LSP sequence in the right subfigure is that shown in Figure 3.
@@
After introducing CODERS, we now investigate its behavior via simulation experiments modeled after data collected by Johnson, (2005). We use the data from this seminal study on careless responding in our empirical application in a later section, where the data are described in more detail.
@@
The survey instrument used by Johnson, (2005) is a 300-item representation of the NEO-PI-R (Goldberg,, 1999) measurement of the Big Five personality traits (openness, conscientiousness, extraversion, agreeableness, and neuroticism; Goldberg,, 1992) by means of five-point Likert scales. Each of the five traits comprises six subdomains (facets) so that a total of $5\times 6=30$ facets are measured. We base our simulation design on the between-item correlation structure and per-item response distributions in the data collected by Johnson, (2005), calculated after removing any presumably careless respondents detected in our empirical application. Figure 5 visualizes the resulting correlation matrix and marginal response distributions. Of the 300 items, 152 are positively scored while 148 are negatively scored. In the correlation matrix, one can identify the five traits and, within each block of within-trait correlations, also the six corresponding facets. Using these correlations and marginal response distributions, we generate rating-scale responses of $n=500$ participants to the $p=300$ items,44endnote: 4We use package simstudy (Goldfeld & Wujciak-Jens,, 2020) to generate the data, for which the target values of correlations only hold approximately on the population level. with the items in the same randomized order as in Johnson, (2005).
@@
Note. The correlation matrix and the marginal response distributions are calculated from the data of Johnson, (2005) after removing presumably careless respondents. Of the $p=300$ items, 152 are positively scored and 148 are negatively scored. The items in both plots are presented in an order such that items measuring the same facet are adjacent.
@@
[…]
@@ Results
We focus on discussing the results for CODERS using both measurements of carelessness (RE and LSP), except when we explicitly state otherwise. Varying the prevalence of careless responding, Figure 6 visualizes the FPR of CODERS in fully attentive respondents for the different carelessness onset regimes (in separate columns). Overall, CODERS rarely flags changepoints in fully attentive respondents. Even with our most liberal choice of significance level $\alpha$ (1%), overrejection increases only slowly to about 3% for 60% prevalence, before it reaches between 3.5% and 5.7% at 80% prevalence (depending on the onset regime). Other than for extremely high prevalence levels, the FPR stays reasonably close to the respective nominal level, indicating that CODERS controls Type I errors fairly well.
@@
In addition, Figure 7 (bottom row) displays the MAE of the estimated onset in partially careless respondents. The prevalence level and onset regime have little effect on the accuracy of the estimates, with an MAE hovering around 2 to 3 depending on the significance level. That is, on average the estimated onset is only two to three items away from the true onset, indicating that CODERS achieves excellent accuracy in estimating carelessness onset.
@@
For random responding (see the first row in Figure 8), CODERS only achieves a decent FNR in the late onset regime with significance level 1% or 0.5% (about 32% and 40%, respectively). In the baseline onset regime, its FNR is higher than 50%. In the early onset regime, even our most liberal significance level of 1% only yields an FNR of about 75%. Nevertheless, this finding should not be mistaken as an intrinsic inability of CODERS to detect random responding. As the results greatly improve the later the onset of carelessness, it seems that the autoencoder needs enough attentive responses to learn the challenging task of distinguishing random responses. In other words, this is an issue of statistical power, which explains why it can be mitigated to some extent by choosing a less conservative significance level. For additional investigation, we apply traditional methods for identifying careless respondents also used by Johnson, (2005). They do not improve upon CODERS (see Appendix C.1), therefore further highlighting the difficulty of detecting random respondents in this simulation design. Coming back to CODERS, it shows similar but far less pronounced behavior for extreme responding (see the second row in Figure 8). The FNR remains decent in the early onset regime, while it is good in the baseline onset regime and excellent (close to 0) in the late onset regime.
To summarize our findings, CODERS performs well for extreme responding and near-perfectly for the two types of invariable responding. On the other hand, it struggles with random responding in this simulation design, where decent power is only achieved in case of late carelessness onset and with a not-too-conservative significance level. However, results for traditional methods (see Appendix C.1) suggest that distinguishing random responding from attentive responding may be an intrinsically difficult problem for the chosen correlation structure and marginal distributions. Overall, we find that (i) CODERS reliably discriminates between attentive and partially careless respondents, as indicated by near-zero false positive rates and low overall false negative rates, (ii) CODERS accurately estimates the location of carelessness onset, and (iii) the performance of CODERS is stable across a wide range of prevalence levels, demonstrating robustness.
@@
[…]
@@
Finally, having so far considered simulations based on personality assessments, we include a simulation inspired by a study from organizational research. This simulated survey is shorter, with $p=198$ items measuring a variety of organizationally relevant constructs, and we apply a different strategy for randomization better suited in organizational contexts. Furthermore, there exist complex relationships between (some of) the measured constructs. The FPR of CODERS is somewhat more elevated for high prevalence levels. Interestingly, the FNR is slightly lower for random responding but slightly higher for extreme responding, resulting in a comparable FNR overall. All in all, the results for CODERS remain relatively similar to those in the simulations based on personality assessments, indicating that its performance may not depend much on the institutional setting of the survey.
@@
After studying the performance of CODERS in simulations inspired by the data from Johnson, (2005), we now re-examine the responses from this seminal study for partial carelessness.
@@
The data collected by Johnson, (2005) concern the five factor model of personality (Goldberg,, 1992), which assumes that variation in a measurement of personality can be explained by the so-called Big Five personality traits (factors): openness, conscientiousness, extraversion, agreeableness, and neuroticism. Each of the Big Five factors can be further decomposed into six subdomains (facets): for instance, depression and anxiety are facets of the neuroticism trait. Consequently, common Big Five instruments, such as the 240-item NEO-PI-R (Costa & McCrae,, 1992), typically measure a total of $5\times 6=30$ facets. The data set of Johnson, (2005) consists of internet-collected responses to a 300-item representation of the NEO-PI-R inventory (Goldberg,, 1999), in which each of the 30 facets is measured by 10 items on a five-point Likert scale.
[…]
In general, it seems likely that survey fatigue is present in a sample of responses to an online questionnaire with a lengthy battery of 300 items. For instance, Bowling et al., (2021) estimate that if one wishes that at least 90% of participants do not respond carelessly to more than 20% of all items, one should not include more than 176 items in an online questionnaire. By applying CODERS—which is specifically designed to identify partially careless responding through a change in response behavior—in addition to the traditional screeners applied by Johnson, (2005), new and complimentary insights into the incidence and nature of (partially) careless responding may emerge.
@@
[…]
@@
For obtaining the autoencoder reconstruction errors in CODERS, we set the number of nodes in the bottleneck layer to 30, i.e., equal the number of facets in the Big Five model. Table 1 summarizes the results of CODERS. Depending on the significance level $\alpha$, CODERS flags between 181 (0.8%) participants for $\alpha=0.001$ and 856 (3.8%) participants for $\alpha=0.01$ as partially careless. On average, the location of the carelessness onset item for flagged participants is estimated around the 145th item, meaning that carelessness onsets on average after nearly half of the 300 items. These findings are broadly in line with the results of Bowling et al., (2021, Figure 1), whose model suggests a low probability that a representative participant starts responding carelessly at the 145th item. The earliest onset occurs at item 15, whereas the latest onset is at item 294. Figure 10 provides a more detailed visual overview of the detected onset items for the flagged participants at different significance levels.
Table 1: Results of CODERS for the data of Johnson, (2005).
@@
[…]
Figure 10: Estimated onset of careless responding for participants flagged by CODERS in the data of Johnson, (2005).
@@
We now turn to analyzing the agreement between CODERS and the two traditional methods applied by Johnson, (2005), personal reliability and psychometric antonym. Following Johnson,, we reverse psychometric antonym scores so that low scores are associated with carelessness in both methods, and we flag respondents as careless whose personal reliability and psychometric antonym scores are lower than $0.3$ and $-0.03$, respectively. The results are shown in Table 2. Personal reliability flags 140 participants (0.6%) as careless, while psychometric antonym flags 284 respondents (1.3%).66endnote: 6We use the R package careless (Yentes & Wilhelm,, 2023) to compute personal reliability and psychometric antonym scores. The use of different software implementations (with potentially different default values) may explain differences with the numbers reported by Johnson, (2005). CODERS, personal reliability, and psychometric antonyms generally show little agreement regarding the detected careless respondents. Personal reliability and psychometric antonym scores correlate by a low value of 0.314, yielding only 11 participants that are flagged by both methods. Depending on the significance level $\alpha$, CODERS and personal reliability have between 2 and 10 flagged participants in common, while the number of jointly flagged respondents for CODERS and psychometric antonym ranges between 9 and 24. All three methods share a single flagged participant.
Table 2: Complementary findings of CODERS and traditional methods for the data of Johnson, (2005).
@@
[…]
@@
Figure 11: Scores and changepoint item detected by CODERS in selected participants from the data of Johnson, (2005).
@@
Table 3: Average scores before and after the changepoint item detected by CODERS in selected participants from the data of Johnson, (2005).
@@
These two cases are representative for the majority of respondents flagged by CODERS, likely starting attentively but starting to respond carelessly at some point. In addition, more respondents flagged by CODERS seem to exhibit inconsistency than invariability, as Johnson, already removed straightliners from the publicly available data set. However, CODERS also identifies participants with remarkably different behavior.
@@
[…]
@@
First, some participants may respond carelessly throughout all survey items (cf. Figure 11(f) and Table 3(f)). Consequently, CODERS is unlikely to flag a changepoint due to an absence of a change in behavior. Nevertheless, the literature has established a plethora of methods for detecting careless respondents. In this sense, CODERS—which is intended for detecting partial carelessness through a change in behavior—is complementary to those existing methods. If one suspects that respondents who have been careless throughout all items are present, we recommend to apply established detection methods following the guidelines of Arthur et al., (2021) and Ward & Meade, (2023). Respondents that are flagged as careless by such methods, but for whom no changepoint was flagged by CODERS, have likely been careless throughout the survey.

### v4

@@ Introduction
We propose to measure response invariability through a novel algorithm inspired by the longstring index of Johnson (2005) that we call LongStringPattern (LSP). This algorithm exploits that invariable careless responses are characterized by content-independent response patterns. A pattern sequence has the defining property that it consists of repeated occurrences of the same response pattern. Consider the sequence of responses $1-2-1-2-\ldots-1-2$, which consists of repeated occurrences of the pattern 121-2. Denoting the length of a pattern by $l$, the above example exhibits a pattern of length $l=2$. Analogously, the sequence $1-2-3-1-2-3-\ldots-1-2-3$ consists of repeated occurrences of the pattern 1231-2-3 of length $l=3$, whereas the straightlining sequence $1-1-\ldots-1$ consists of repeated occurrences of 1, which is a pattern of length $l=1$.
Define $l\textsf{-pattern}_{ij}$ as the number of consecutive items in the longest sequence of a recurring pattern of length $l$ in which the response of participant $i$ to item $j$ takes part, or assign the value 1 if the response is not part of a recurring pattern of length $l$. Consider the following two illustrative examples. First, for $l=2$, the response sequence 121214354541-2-1-2-1-4-3-5-4-5-4 will be assigned the 2-pattern sequence 444411144444-4-4-4-1-1-1-4-4-4-4 because the first four responses and last four responses comprise two occurrences of the distinct patterns 121-2 and 545-4, respectively, which are both of length $l=2$. The central three responses (1431-4-3) are not part of any recurring pattern of length $l=2$, hence they are each assigned the value 1. Second, for a recurring pattern of length $l=1$ (i.e., consecutive identical responses), the response sequence 3233141113-2-3-3-1-4-1-1-1 is assigned the 1-pattern sequence 1122113331-1-2-2-1-1-3-3-3, since the subsequences of consecutive identical responses 333-3 and 1111-1-1 comprise two and three responses, respectively. [^3] The 1-pattern sequence can be seen as a generalization of the the longstring index of Johnson (2005), which we recover by calculating $\max_{1\leq j\leq p}1\textsf{-pattern}_{ij}$ for each respondent $i=1,\dots,n$. In short, higher values in the $l$-pattern sequence capture a tendency to respond in fixed patterns of length $l$. If a participant starts engaging in invariable careless responding, we expect a changepoint in the $l$-pattern sequence (for some $l$) from relatively low to relatively high values, assuming that the survey design precludes such patterns in attentive responses.
@@
for participant $i=1,\dots,n$ and item $j=1,\dots,p$. We call the series of within-participant values $\textsf{LSP}_{ij}$ across all items a LongStringPattern (LSP) sequence. In this paper, we vary $l$ from 1 to the number of answer categories in the survey. The rationale behind this choice is that a careless respondent may repeat, e.g., the pattern 123451-2-3-4-5 in case of five answer categories. At the same time, we consider it unlikely that careless respondents choose complicated individual patterns whose length exceeds the number of answer categories.
We provide a detailed description of the LSP algorithm in Appendix B. An example of an LSP sequence is shown in Figure 3, which shows simulated responses of a hypothetical participant in a 300-item questionnaire (5 answer categories per item) who starts responding carelessly using the recurring pattern 131-3 from the 245th item, resulting in systematically larger LSP values from that item onward.
@@
Note. The simulated participant started responding carelessly with the recurring pattern 131-3 from the 245th item onward.
@@
As motivated in the previous two sections, we expect a changepoint in any of the measurements of carelessness from the onset of careless responding (cf. Figures 2 and 3), and no changepoint in the absence of carelessness. We therefore propose to apply the nonparametric cumulative sum self-normalization test of Shao and Zhang (2010), which was developed in the context of time series analysis. It searches for the location of a possible changepoint in a multivariate series and has attractive theoretical guarantees (see Shao and Zhang, 2010; Zhao et al., 2022; and our detailed description in Appendix B).
For each point in a multivariate series (we use measurements of carelessness associated with each questionnaire item), this changepoint detection test calculates a statistic that compares a quantity of interest (we take the mean) in the periods before and after the current point (item). If the maximum of this statistic exceeds a critical value, the point (item) where this maximum occurs is flagged as a changepoint. Otherwise no changepoint is flagged. The critical value thereby depends only on a pre-specified significance level $\alpha$ and the dimension of the multivariate series. Importantly, it controls the size of the test regardless of the distribution of the individual series. This gives CODERS great flexibility, as it allows us to use any combination of measurements of carelessness without making possibly unrealistic assumptions on their distributions or obtaining cutoff values in an ad-hoc manner. In line with literature that suggests that there should be overwhelming evidence when labeling respondents as careless (e.g., Huang et al., 2012), we recommend using a conservative significance level no greater than 1% for identifying changepoints.
@@
Note. The left subfigure shows the output from CODERS for a simulated partially inconsistent respondent who starts responding randomly from the 183rd item onward, while the right subfigure does so for a simulated partially invariable respondent who starts repeating the pattern 131-3 from the 245th item onward. Each subfigure visualizes the autoencoder reconstruction errors (RE), LongStringPattern (LSP) sequence, and test statistics associated with the corresponding participant’s responses. The blue solid vertical lines indicate the carelessness onset items identified by CODERS at significance level 0.1%, which correspond to item 181 (two items before true onset) for the partially inconsistent respondent, and to the true onset item 245 for the partially invariable respondent. The dashed horizontal line is the critical value of the test statistic at the 0.1% significance level (192.5; see Table B2 in Appendix B). The reconstruction errors in the left subfigure are those shown in Figure 2, and the LSP sequence in the right subfigure is that shown in Figure 3.
As an illustrative example, consider Figure 4(a). It illustrates two dimensions computed from a given participant’s responses, namely autoencoder reconstruction errors and the LSP sequence. The former series has been shown before in Figure 2. The bottom panel of Figure 4(a) shows the test statistics computed from this two-dimensional series, where the maximum value at the 181st item clearly exceeds the critical value. That is, CODERS identifies inconsistent carelessness from the 181st item onward, as the reconstruction errors are in general much larger than before. Indeed, this simulated participant started choosing responses at random just two items after the estimated onset. Figure 4(b) is similar, but computed from a different respondent, whose LSP sequence has been shown before in Figure 3. CODERS flags a changepoint at the 245th item. From this item onward, the LSP values are clearly elevated, indicating invariable carelessness. In fact, this simulated participant started repeating the pattern 131-3 at the identified item.
@@
After introducing CODERS, we now investigate its behavior via simulation experiments modeled after data collected by Johnson (2005). We use the data from this seminal study on careless responding in our empirical application in a later section, where the data are described in more detail. Unfortunately, it is not possible to compare CODERS with the approach of Yu and Cheng (2019), which is the only other method for detecting the onset of careless responding that we are aware of, as no R code seems to be available.
@@
The survey instrument used by Johnson (2005) is a 300-item representation of the NEO-PI-R (Goldberg, 1999) measurement of the Big Five personality traits (openness, conscientiousness, extraversion, agreeableness, and neuroticism; Goldberg, 1992) by means of five-point Likert scales. Each of the five traits comprises six subdomains (facets) so that a total of $5\times 6=30$ facets are measured. We base our simulation design on the between-item correlation structure and per-item response distributions in the data collected by Johnson (2005), calculated after removing any presumably careless respondents detected in our empirical application. Figure 5 visualizes the resulting correlation matrix and marginal response distributions. Of the 300 items, 152 are positively scored while 148 are negatively scored. In the correlation matrix, one can identify the five traits and, within each block of within-trait correlations, also the six corresponding facets. Using these correlations and marginal response distributions, we generate rating-scale responses of $n=500$ participants to the $p=300$ items, [^4] We use package simstudy (Goldfeld and Wujciak-Jens, 2020) to generate the data, for which the target values of correlations only hold approximately on the population level. with the items in the same randomized order as in Johnson (2005).
@@
Note. The correlation matrix and the marginal response distributions are calculated from the data of Johnson (2005) after removing presumably careless respondents. Of the $p=300$ items, 152 are positively scored and 148 are negatively scored. The items in both plots are presented in an order such that items measuring the same facet are adjacent.
@@
From the carelessness onset item onward, we generate the responses from one of the following four careless response types: (i) random responding (choosing answer categories completely at random), (ii) extreme responding (randomly choosing between the two most extreme answer categories), (iii) straightlining (repeating the same, randomly determined answer category), or (iv) pattern responding (repeating a fixed pattern such as 1231-2-3 or 545-4, which is randomly determined from all possible patterns of length two to five). Specifically, we assign these careless response types to the partially careless respondents so that there are $\lfloor\gamma n/4\rfloor$ respondents for each response type, where $\lfloor\ .\ \rfloor$ denotes the operator for rounding down to the nearest integer.
@@ Results
We focus on discussing the results for CODERS using both measurements of carelessness (RE and LSP), except when we explicitly state otherwise. Varying the prevalence of careless responding, Figure 6 visualizes the FPR of CODERS in fully attentive respondents for the different carelessness onset regimes (in separate columns). Overall, CODERS rarely flags changepoints in fully attentive respondents. Even with our most liberal choice of significance level $\alpha$ (1%), overrejection increases only slowly to about 3% for 60% prevalence, before it reaches between 3.0% and 5.6% at 80% prevalence (depending on the onset regime). Other than for extremely high prevalence levels, the FPR stays reasonably close to the respective nominal level, indicating that CODERS controls Type I errors fairly well.
@@
In addition, Figure 7 (bottom row) displays the MAE of the estimated onset in partially careless respondents. The prevalence level and onset regime have little effect on the accuracy of the estimates, with an MAE hovering around 2 to 3, depending on the significance level. That is, on average the estimated onset is only two to three items away from the true onset, indicating that CODERS achieves excellent accuracy in estimating carelessness onset.
@@
For random responding (see the first row in Figure 8), CODERS only achieves a decent FNR in the late onset regime with significance level 1% or 0.5% (about 32% and 40%, respectively). In the baseline onset regime, its FNR is higher than 50%. In the early onset regime, even our most liberal significance level of 1% only yields an FNR of about 75%. Nevertheless, this finding should not be mistaken as an intrinsic inability of CODERS to detect random responding. As the results greatly improve the later the onset of carelessness, it seems that the autoencoder needs enough attentive responses to learn the challenging task of distinguishing random responses. In other words, this is an issue of statistical power, which explains why it can be mitigated to some extent by choosing a less conservative significance level. For additional investigation, we apply traditional methods for identifying careless respondents also used by Johnson (2005). They do not improve upon CODERS (see Appendix C), therefore further highlighting the difficulty of detecting random respondents in this simulation design. Coming back to CODERS, it shows similar but far less pronounced behavior for extreme responding (see the second row in Figure 8). The FNR remains decent in the early onset regime, while it is good in the baseline onset regime and excellent (close to 0) in the late onset regime.
To summarize our findings, CODERS performs well for extreme responding and near-perfectly for the two types of invariable responding. On the other hand, it struggles with random responding in this simulation design, where decent power is only achieved in case of late carelessness onset and with a not-too-conservative significance level. However, results for traditional methods (see Appendix C) suggest that distinguishing random responding from attentive responding may be an intrinsically difficult problem for the chosen correlation structure and marginal distributions. Overall, we find that (i) CODERS reliably discriminates between attentive and partially careless respondents, as indicated by near-zero false positive rates and low overall false negative rates, (ii) CODERS accurately estimates the location of carelessness onset, and (iii) the performance of CODERS is stable across a wide range of prevalence levels, demonstrating robustness.
@@
Second, we vary the number of participants by setting $n=248$ and $n=\text{1,000}$. [^5] Since we generate four types of partially careless respondents in the simulations, the sample size should be divisible by 4, explaining the unusual choice of $n=248$ (rather than $n=250$). The overall FNR in partially careless respondents slightly increases for the smaller sample size and slightly decreases for the larger sample size, with bigger changes being observed for random respondents. That is, the effect of the sample size on power is relatively small except for random responding, where further improvements in performance may be possible for even larger sample sizes.
@@
After studying the performance of CODERS in simulations inspired by the data from Johnson (2005), we now re-examine the responses from this seminal study for partial carelessness.
@@
The data collected by Johnson (2005) concern the five factor model of personality (Goldberg, 1992), which assumes that variation in a measurement of personality can be explained by the so-called Big Five personality traits (factors): openness, conscientiousness, extraversion, agreeableness, and neuroticism. Each of the Big Five factors can be further decomposed into six subdomains (facets): for instance, depression and anxiety are facets of the neuroticism trait. Consequently, common Big Five instruments, such as the 240-item NEO-PI-R (Costa and McCrae, 1992), typically measure a total of $5\times 6=30$ facets. The data set of Johnson (2005) consists of internet-collected responses to a 300-item representation of the NEO-PI-R inventory (Goldberg, 1999), in which each of the 30 facets is measured by 10 items on a five-point Likert scale.
Johnson (2005) identifies straightliners via the longstring index (which the author proposed in that paper) and inconsistent careless respondents via personal reliability (Jackson, 1976) and psychometric antonym (e.g., Meade and Craig, 2012) scores. The data from this study are made publicly available by the author at https://osf.io/sxeq5/. Importantly, from the original sample of 23,994 participants, the author removed 918 participants that were likely duplicates, 1,455 participants with too many straightlining responses, and 688 participants with too many missing responses, leaving $n=\text{20,933}$ participants in the provided data set. Unfortunately, the raw data is not available at the above link. The minimal preprocessing that we applied to the provided data set is described in Appendix D.
In general, it seems likely that survey fatigue is present in a sample of responses to an online questionnaire with a lengthy battery of 300 items. For instance, Bowling et al. (2021) estimate that if one wishes that at least 90% of participants do not respond carelessly to more than 20% of all items, one should not include more than 176 items in an online questionnaire. By applying CODERS—which is specifically designed to identify partially careless responding through a change in response behavior—in addition to the traditional screeners applied by Johnson (2005), new and complimentary insights into the incidence and nature of (partially) careless responding may emerge.
@@
Throughout the discussion of the results, it is important to keep in mind that Johnson (2005) already removed straightliners from the publicly available data set. To maintain consistency in reporting, any percentages of participants refer to a baseline sample size of $N=\text{22,448}$ (i.e., including straightliners). For instance, the 1,455 straightliners found by Johnson (2005) amount to 6.5% of participants.
@@
For obtaining the autoencoder reconstruction errors in CODERS, we set the number of nodes in the bottleneck layer to 30, i.e., equal the number of facets in the Big Five model. Table 1 summarizes the results of CODERS. Depending on the significance level $\alpha$, CODERS flags between 181 (0.8%) participants for $\alpha=0.001$ and 856 (3.8%) participants for $\alpha=0.01$ as partially careless. On average, the location of the carelessness onset item for flagged participants is estimated around the 145th item, meaning that carelessness onsets on average after nearly half of the 300 items. The earliest onset occurs at item 15, whereas the latest onset is at item 294. Figure 10 provides a more detailed visual overview of the detected onset items for the flagged participants at different significance levels.
Table 1: Results of CODERS for the data of Johnson (2005).
@@
Note. The methods are applied to the responses of $n=\text{20,993}$ participants to $p=300$ items of a Big Five personality measurement. Among others, participants with too many straightlining responses have already been removed from the data provided by Johnson. For consistent reporting, percentages refer to a baseline sample size of $N=\text{22,448}$, which includes straightliners.
Figure 10: Estimated onset of careless responding for participants flagged by CODERS in the data of Johnson (2005).
@@
Table 2: Complementary findings of CODERS and traditional methods for the data of Johnson (2005).
@@
Note. The methods are applied to the responses of $n=\text{20,993}$ participants to $p=300$ items of a Big Five personality measurement. Among others, participants with too many straightlining responses have already been removed from the data provided by Johnson. For consistent reporting, percentages refer to a baseline sample size of $N=\text{22,448}$, which includes straightliners. Too small percentages are omitted for better readability.
We now turn to analyzing the agreement between CODERS and the two traditional methods applied by Johnson (2005), personal reliability and psychometric antonym. Following Johnson, we reverse psychometric antonym scores so that low scores are associated with carelessness in both methods, and we flag respondents as careless whose personal reliability and psychometric antonym scores are lower than $0.3$ and 0.03-0.03, respectively. The results are shown in Table 2. Personal reliability flags 140 participants (0.6%) as careless, while psychometric antonym flags 284 respondents (1.3%). [^6] We use the R package careless (Yentes and Wilhelm, 2023) to compute personal reliability and psychometric antonym scores. The use of different software implementations (with potentially different default values) may explain differences with the numbers reported by Johnson (2005). CODERS, personal reliability, and psychometric antonyms generally show little agreement regarding the detected careless respondents. Personal reliability and psychometric antonym scores correlate by a low value of 0.314, yielding only 11 participants that are flagged by both methods. Depending on the significance level $\alpha$, CODERS and personal reliability have between 2 and 10 flagged participants in common, while the number of jointly flagged respondents for CODERS and psychometric antonym ranges between 9 and 24. All three methods share a single flagged participant.
@@
Figure 11: Scores and changepoint item detected by CODERS in selected participants from the data of Johnson (2005).
@@
Table 3: Average scores before and after the changepoint item detected by CODERS in selected participants from the data of Johnson (2005).
@@
These two cases are representative for the majority of respondents flagged by CODERS, likely starting attentively but starting to respond carelessly at some point. In addition, more respondents flagged by CODERS seem to exhibit inconsistency than invariability, as Johnson already removed straightliners from the publicly available data set. However, CODERS also identifies participants with remarkably different behavior.
@@
Finally, the example in Figure 11(f) and Table 3(f) corresponds to a participant flagged by psychometric antonym but not CODERS. A visual inspection confirms that there is indeed no clear changepoint in either the reconstruction errors or the LSP sequence. However, the reconstruction errors are similar in magnitude and variability to those found in inconsistent response periods from other examples. To investigate this further, we computed average reconstruction errors across all $p=300$ items for the $n=\text{20,993}$ participants. Only 6 participants have a higher average reconstruction error (cf. Figure D1 in Appendix D), providing overwhelming evidence that this participant may be an inconsistent careless respondent throughout the whole questionnaire.
@@
First, some participants may respond carelessly throughout all survey items (cf. Figure 11(f) and Table 3(f)). Consequently, CODERS is unlikely to flag a changepoint due to an absence of a change in behavior. Nevertheless, the literature has established a plethora of methods for detecting careless respondents. In this sense, CODERS—which is intended for detecting partial carelessness through a change in behavior—is complementary to those existing methods. If one suspects that respondents who have been careless throughout all items are present, we recommend to apply established detection methods following the guidelines of Arthur et al. (2021) and Ward and Meade (2023). Respondents that are flagged as careless by such methods, but for whom no changepoint was flagged by CODERS, have likely been careless throughout the survey.